OpenStudy (lgbasallote):

\(\LARGE \int \frac{4x+3}{\sqrt{5 + 9x - x^2 }} dx\) is this completing the square then inverse trig?

OpenStudy (lgbasallote):

here it is @Mimi_x3 :D

OpenStudy (mimi_x3):

looks like complete the square then a u-sub

OpenStudy (mimi_x3):

wait..let me try it first..

OpenStudy (lgbasallote):

okay then :D though completing the square seems tough since itll be (9/2 -x)^2 which looks complicated

OpenStudy (mimi_x3):

well, have you completed the square first

OpenStudy (lgbasallote):

do you mean in this problem? hmm like i said it seemed complicated

OpenStudy (lgbasallote):

i'll have \(\LARGE \sqrt{(\frac{9}{2} - x)^2 -\frac{61}{4}}\) lol

OpenStudy (mimi_x3):

wolfram alpha says its.. \[\frac{101}{4}-\left(x-\frac{9}{2}\right)^{2} \] too tired to do it..lol

OpenStudy (lgbasallote):

o.O

OpenStudy (lgbasallote):

well then assuming that's it...then.. this is arcsin right...

OpenStudy (lgbasallote):

how do i get rid of the numerator? one of my common problems in these types

OpenStudy (mimi_x3):

You did something wrong when you completed the sqaue. \[\left(-x+\frac{9}{2}\right) ^{2} -\left(\frac{9}{2}\right) ^{2} +5\] then.. \[\frac{101}{4}-\left(x-\frac{9}{2}\right)^{2} \]

OpenStudy (mimi_x3):

Then you have to u-sub

OpenStudy (lgbasallote):

i dont see the u sub lol

OpenStudy (mimi_x3):

\[\large \frac{4x+3}{\sqrt{(-x+\frac{9}{2})^{2}-\frac{101}{4}}} \]

OpenStudy (mimi_x3):

hmm..

OpenStudy (mimi_x3):

wiat

OpenStudy (mimi_x3):

breaking it up into two fractions might work

OpenStudy (mimi_x3):

brb 5 min

OpenStudy (lgbasallote):

\(\LARGE \int \frac{4x}{\sqrt{(-x + \frac{9}{2})^2 - \frac{101}{4}}} + \frac{3}{\sqrt{(-x + \frac{9}{2})^2 - \frac{101}{4}}}\)

OpenStudy (lgbasallote):

if u = -x + \(\frac{9}{2}\) then du = -dx that does not satisfy 4x:P

OpenStudy (mimi_x3):

will work for the second one i will have to think about the first

OpenStudy (lgbasallote):

for the second one it will be a trig sub right?

OpenStudy (mimi_x3):

its going to get into a big mess..i think that its a trig sub

OpenStudy (mimi_x3):

i think the first one is a trig sub

OpenStudy (lgbasallote):

lol this is complicated =))

OpenStudy (mimi_x3):

no; the second one is easy.. man i can't think..too tired..i think there is a better way..

OpenStudy (lgbasallote):

this looks so complicated to me :/ yet i cannot see any other way

OpenStudy (mimi_x3):

hmph; i give up im going to sleep..will think of a better way later if no one answers it..

OpenStudy (lgbasallote):

okay thanks for the help mimi :)

OpenStudy (zarkon):

why are you using \[\sqrt{(-x+\frac{9}{2})^{2}-\frac{101}{4}}\] as the bottom of your integrand?

OpenStudy (lgbasallote):

to make it look like an arcsin?

OpenStudy (lgbasallote):

i mean the one with 1/a arcsin u/a

OpenStudy (zarkon):

but when you complete the square you get \[\frac{101}{4}-\left(x-\frac{9}{2}\right)^{2}\]

OpenStudy (anonymous):

\[\frac{2(2x-9) + 18 + 3}{\sqrt{5 + 9x - x^2}} = 2\frac{2x - 9}{\sqrt{5 + 9x - x^2}} + \frac{21}{\sqrt{5 + 9x - x^2}}\] \[5 + 9x - x^2=t^2 \implies 2t \frac{dt}{dx} =9 -2x \] Perform the t substitution in the first one and complete squares in the second one. It should work.

OpenStudy (zarkon):

don't you have it backwards

OpenStudy (lgbasallote):

what did ishaan do the last step

OpenStudy (anonymous):

Hmm? lgb? I tried to cancel (get rid of the x term) the numerator by the substitution of the polynomial under square root.

OpenStudy (lgbasallote):

i meant this 5+9x−x^2=t^2

OpenStudy (anonymous):

Basic substitution. I didn't want to have square roots.

OpenStudy (anonymous):

after the substitution.*

OpenStudy (lgbasallote):

oh i see...well how you gonna plug in 9 - 2x:P

OpenStudy (anonymous):

\[2\frac{2x - 9}{\sqrt{5 + 9x - x^2}} dx = 2\frac{2x - 9}{\sqrt{5 + 9x - x^2}} \cdot \frac{2t}{9-2x}\cdot dt = 2\cdot \frac{-(9-2x)}{t}\cdot \frac{2t}{9-2x}dt\]